OpenStudy (joppei):

Can someone tell me how the RMSD (root mean square devition) differs from the standard deviation (which I am familiar with; the formula's look so alike...)

OpenStudy (amistre64):

RMSD is the sqrt of the mean of the sum of the squares \[\sqrt{\frac{x_1^2+x_2^2+x_3^2+...+x_n^2}{n}}\] standard deviation is the sqrt of variance \[\sqrt{\frac{(x_1-\mu)^2+(x_2-\mu)^2+(x_3-\mu)^2+...+(x_n-\mu)^2}{n}}\]

OpenStudy (joppei):

If i did it correctly the RMSD of these numbers: 3, 4, 5, 1, 1, 2 is 3.06? while the mean is 2.67 and the standard deviation is 1.33. What does the RMSD tell me?

OpenStudy (amistre64):

not too sure what it tells us, other than its some other attempt to define spread or density of the data set

OpenStudy (joppei):

maybe it can be used to relate the deviation between individual values in stead of relating them all the the mean. So if you pick two numbers from my sequence they will be 3.06 seperated from eachother.. that number seems rather large

OpenStudy (amistre64):

the standard deviation relates the data set to a normal distribution with a mean of zero. If you were to put the mean into the set, and subtract the mean from all the data points .... youd simply be zeroing out the mean. And of course using smaller values in the operation. the RMS doesnt move the set and therefore uses larger values in the process. the spread itself remains the same for each data set .... my lack of experience as a whole with the RMS stuff doesnt allow me to see what, if any, significance it would play tho

OpenStudy (amistre64):

im also coming across differening material as to how to calculate the RMS the one I presented above was the Wolframs version of it; but i see others of it that suggest that the sample mean, instead of the population mean, is used in the operation :/

OpenStudy (joppei):

statisticians huh, the one i am supposed to understand is used in molecular modeling and the formula is:

OpenStudy (joppei):

k is the position of a particle

OpenStudy (joppei):

no sorry it is just a particle, not a position

OpenStudy (amistre64):

that looks like the mean of the squares, of the differences between all the pairs of observations

OpenStudy (joppei):

the root of those squares

OpenStudy (joppei):

its like an unscaled standard deviation

OpenStudy (amistre64):

3-{3, 4, 5, 1, 1, 2): {0,1,2,2,2,1} 4-{3, 4, 5, 1, 1, 2): {1,0,1,3,3,2} 5-{3, 4, 5, 1, 1, 2): {2,1,0,4,4,3} 1-{3, 4, 5, 1, 1, 2): {2,3,4,0,0,1} 1-{3, 4, 5, 1, 1, 2): {2,3,4,0,0,1} 2-{3, 4, 5, 1, 1, 2): {1,2,3,1,1,0} {0,1,2,2,2,1} sum: 0 1 4 4 4 1 = 14 {1,0,1,3,3,2} 1 0 1 9 9 4 = 24 {2,1,0,4,4,3} 4 1 0 16 16 9 = 46 {2,3,4,0,0,1} {2,3,4,0,0,1} 4 9 16 0 0 1 = 30 30 {1,2,3,1,1,0} 1 4 9 1 1 0 = 16 -------- 160 , for a sum ?

OpenStudy (amistre64):

whats the dividor in your set up, all the squares? or the number of the original set?

OpenStudy (joppei):

it is multiplied by the reciproke of n

OpenStudy (joppei):

but also you have to root it first right?

OpenStudy (amistre64):

yes, if my text is good, then sqrt(160/6^2); or about 2.11

OpenStudy (joppei):

which is smaller than the standard deviation.. but it seems about right i think

OpenStudy (joppei):

i think this might be correct

OpenStudy (amistre64):

theres a setup on one of the pages i see:\[2\sigma^2=\frac{\sum(X_i-X_j)^2}{N^2}\] \[\sigma^2=\frac{\sum(X_i-X_j)^2}{2N^2}\] \[\sigma=\sqrt{\frac{\sum(X_i-X_j)^2}{2N^2}}\]

OpenStudy (joppei):

and it gives the average difference between two molecules from the model at a certain timepoint if i'm not mistaking

OpenStudy (amistre64):

lol, that last one didnt seem to go over to well withthe wolf

OpenStudy (amistre64):

but yes, the average distance between all the observed pairs

OpenStudy (joppei):

so in my case i would have 6 numbers so 7 molecules.. you mean wolfram mathematica? :P

OpenStudy (amistre64):

yes, the wolf :)

OpenStudy (joppei):

=D its stilla bit hazy but i think i can work with this. sorry for taking so much of your time and thnx!

OpenStudy (amistre64):

it was fun :) good luck and all